# Detecting High-Stakes Interactions with Activation Probes

### v3

@@ Abstract
Monitoring is an important aspect of safely deploying Large Language Models (LLMs). This paper examines activation probes for detecting “high-stakes” interactions—where the text indicates that the interaction might lead to significant harm—as a critical, yet underexplored, target for such monitoring. We evaluate several probe architectures trained on synthetic data, and find them to exhibit robust generalization to diverse, out-of-distribution, real-world data. Probes’ performance is comparable to that of prompted or finetuned medium-sized LLM monitors, while offering computational savings of six orders-of-magnitude. These savings are enabled by reusing activations of the model that is being monitored. Our experiments also highlight the potential of building resource-aware hierarchical monitoring systems, where probes serve as an efficient initial filter and flag cases for more expensive downstream analysis.
We release our novel synthetic dataset and the codebase at https://github.com/arrrlex/models-under-pressure.
@@ 2.1 Probes
Let $A=[\bm{a}_{1},\dots,\bm{a}_{S}]\in\mathbb{R}^{S\times D}$ be the residual activations before the layer norm of some arbitrary layer in the LLM, where $S$ denotes the sequence length, and $D$ the residual dimension. The output of a probe is given by $p_{\theta}(A):=\sigma(f_{\theta}(A))$, where $\sigma$ is the sigmoid function and $\theta$ is a vector (or set of vectors) of learned parameters. $f_{\theta}$ is dependent on the chosen probe architecture, which includes design decisions such as how to aggregate the activations across the sequence dimension. We considered taking the mean, the maximum, the last token score, an average weighted by probe score (“Softmax Probe”), an average weighted by a learned linear function (“Attention Probe”), and the maximum score after taking a rolling-window average (“Max of Rolling Means”). To the best of our knowledge, Softmax and Max of Rolling Means have not been used for activation probing before. The Attention probe was introduced by Kantamneni et al. (2025). We trained all probes using the AdamW optimizer (Loshchilov & Hutter, 2019). Details, including hyperparameters and mathematical definitions, can be found in Section˜A.1.
@@ 2.2 Baselines
Following standard practice in interpretability (Bhalla et al., 2025; Wu et al., 2025), we compared our white-box probes against two black-box text classification methods: finetuning (Inan et al., 2023a; Zeng et al., 2025; Mao et al., 2023; Howard & Ruder, 2018; Bucher & Martini, 2024; Zhao et al., 2024; Wei et al., 2023; Pangakis & Wolken, 2024) and prompting. We implemented both types of baselines for various LLMs of the Llama-3 (Grattafiori et al., 2024) and Gemma-3 (Team et al., 2025) model families.
Finetuned baselines. Following standard practice (Sanh et al., 2020; Arora et al., 2022) we converted autoregressive LLMs into classifiers by removing the final unembedding layer and attaching a linear classification head.
We then finetuned the whole model on our synthetic training dataset. For further details, see Section˜A.2.
@@ 2.4 Training and Evaluation Datasets
To ensure unambiguous labels, we filtered samples using GPT-4o-assigned stakes scores (1–10) and confidence scores (1–10), removing ambiguous stakes (scores 4–7) or low-confidence samples (<8). We also applied a bag-of-words classifier to eliminate samples dominated by confounding words (i.e. words such as “minor” or “trivial” whose presence overly influenced bag-of-words classification).
Specific details on training data generation and filtering criteria are elaborated in Appendix˜C.
@@ 3 Experiments
Using our synthetic training data (Section˜2.4), we trained probes with various architectures (Section˜2.1) and baseline methods (Section˜2.2) to explore how well different techniques detect the high-stakes contexts in our evaluation datasets (Section˜2.4).
Our probes are trained on the residual-stream activations of a single layer of Llama-3.3-70B (Grattafiori et al., 2024). This layer—31 of 80—was chosen via cross-validation using an early probe design (see Section˜A.1). Additional experiments on smaller models, such as Llama-3.1-8B and Gemma-3-27B, can be found in Section˜B.7. As training probes is resource-efficient and labels of reasonable quality can realistically be generated using GPT-4o or similar models, we also study how much training on deployment samples (from the development splits, in our case) can help with better capturing the monitored concept.
@@ 5 Related Work
Linear probes. Linear probes (Alain & Bengio, 2018) analyze neural network representations and have been used to understand model internals (Belinkov & Glass, 2019; Tenney et al., 2019; Jin et al., 2025). Alternative designs include Linear Artificial Tomography (Zou et al., 2023) and Mass-Means probes (Marks & Tegmark, 2023). Our work demonstrates how probes can act as computationally effective initial filters. Sparse Auto-Encoders (SAEs) (Cunningham et al., 2023; Bricken et al., 2023) offer an alternative approach via unsupervised clustering, but recent work suggests reliability issues for downstream monitoring tasks (Chanin et al., 2024; Thread, 2023; Kantamneni et al., 2025). We therefore focus on activation probes.
@@
Monitoring for safety-relevant properties. Numerous works detect undesirable LLM behavior via finetuning (Inan et al., 2023a; Zeng et al., 2024). Linear probes offer a cheaper alternative for investigating safety-related concepts like truth and honesty (Marks & Tegmark, 2023; Azaria & Mitchell, 2023; Li et al., 2022; Burns et al., 2023; Panickssery et al., 2023), outperforming prompting whether a statement is true on smaller models in-distribution Liu et al. (2023). While some doubt remains about generalization—Levinstein & Herrmann (2024) show probes from Azaria & Mitchell (2023) fail simple generalization tests, and Belinkov (2021) argues for general skepticism—more recent studies demonstrate robust generalization across transforms and distribution shifts (Wagner, 2024; Roger, 2023). Most recently, Goldowsky-Dill et al. (2025) show deception can be reliably detected across diverse datasets. We apply linear probes to the novel setting of high-stakes user interactions, with rigorous benchmarking against LLM monitors.
Several concurrent works explore related monitoring approaches with different technical choices. Tillman & Mossing (2025) use logistic regression and SAE-based probes on GPT-4o, finding that some probes outperform prompting. This contrasts with our finding that Llama-3.3-70B probes require deployment samples to exceed prompting baselines. We attribute this to: (1) our synthetic-to-real-world OOD evaluation versus their real-data training with smaller distributional shifts, and (2) our extensively optimized prompting baseline (Section˜A.2). Cunningham et al. (2025) highlight the cost-effectiveness of reusing representations for monitoring using either linear probes or partially fine-tuned models, and also explore such monitors as an initial stage in a hierarchical setup. While our work focuses on high-stakes scenarios, theirs centers on jailbreak and harmfulness detection, with many findings aligning with and supporting our own. Oldfield et al. (2025) introduce Truncated Polynomial Classifiers, which augment linear probes with higher-order feature interactions to allow a small amount of dynamic compute scaling within a single activation monitor model.
Decision-making in High-Stakes situations. Human responses under high-stakes conditions have been extensively studied in psychology. Kunreuther et al. (2002) shows how high-stakes scenarios lead to suboptimal decision making, extended by Sosnowski & Brosnan (2023) to nonhuman species. In AI settings, Ren et al. (2025) demonstrates that pressured prompts incentivize LLMs toward dishonesty, while Motwani et al. (2024) shows high-stakes contexts can induce unintended collusion and illicit behaviors like insider trading.

### v4

@@ Abstract
Monitoring is an important aspect of safely deploying Large Language Models (LLMs). This paper examines activation probes for detecting “high-stakes” interactions—where the text indicates that the interaction might lead to significant harm—as a critical, yet underexplored, target for such monitoring. We evaluate several probe architectures trained on synthetic data, and find them to exhibit robust generalization to diverse, out-of-distribution, real-world data. Probes’ performance is comparable to that of prompted or finetuned medium-sized LLM monitors, while offering computational savings of six orders-of-magnitude. These savings are enabled by reusing activations of the model that is being monitored. Our experiments also highlight the potential of building resource-aware hierarchical monitoring systems, where probes serve as an efficient initial filter and flag cases for more expensive downstream analysis. We release our novel synthetic dataset and the codebase at https://github.com/arrrlex/models-under-pressure.
@@ 2.1 Probes
Let $A=[\bm{a}_{1},\dots,\bm{a}_{S}]\in\mathbb{R}^{S\times D}$ be the residual activations before the layer norm of some arbitrary layer in the LLM, where $S$ denotes the sequence length, and $D$ the residual dimension. The output of a probe is given by $p_{\theta}(A):=\sigma(f_{\theta}(A))$, where $\sigma$ is the sigmoid function and $\theta$ is a vector (or set of vectors) of learned parameters. $f_{\theta}$ is dependent on the chosen probe architecture, which includes design decisions such as how to aggregate the activations across the sequence dimension. We considered taking the mean, the maximum, the last token score, an average weighted by probe score (“Softmax Probe”), an average weighted by a learned linear function (“Attention Probe”), and the maximum score after taking a rolling-window average (“Max of Rolling Means”). To the best of our knowledge, Softmax and Max of Rolling Means have not been used for activation probing before. The Attention probe was introduced by Kantamneni et al. (2025). We trained all probes using the AdamW optimizer (Loshchilov and Hutter, 2019). Details, including hyperparameters and mathematical definitions, can be found in Section˜A.1.
@@ 2.2 Baselines
Following standard practice in interpretability (Bhalla et al., 2025; Wu et al., 2025), we compared our white-box probes against two black-box text classification methods: finetuning (Inan et al., 2023a; Zeng et al., 2025; Mao et al., 2023; Howard and Ruder, 2018; Bucher and Martini, 2024; Zhao et al., 2024; Wei et al., 2023; Pangakis and Wolken, 2024) and prompting. We implemented both types of baselines for various LLMs of the Llama-3 (Grattafiori and others, 2024) and Gemma-3 (Team and others, 2025) model families.
Finetuned baselines. Following standard practice (Sanh et al., 2020; Arora et al., 2022) we converted autoregressive LLMs into classifiers by removing the final unembedding layer and attaching a linear classification head. We then finetuned the whole model on our synthetic training dataset. For further details, see Section˜A.2.
@@ 2.4 Training and Evaluation Datasets
To ensure unambiguous labels, we filtered samples using GPT-4o-assigned stakes scores (1–10) and confidence scores (1–10), removing ambiguous stakes (scores 4–7) or low-confidence samples (<8). We also applied a bag-of-words classifier to eliminate samples dominated by confounding words (i.e. words such as “minor” or “trivial” whose presence overly influenced bag-of-words classification). Specific details on training data generation and filtering criteria are elaborated in Appendix˜C.
@@ 3 Experiments
Using our synthetic training data (Section˜2.4), we trained probes with various architectures (Section˜2.1) and baseline methods (Section˜2.2) to explore how well different techniques detect the high-stakes contexts in our evaluation datasets (Section˜2.4). Our probes are trained on the residual-stream activations of a single layer of Llama-3.3-70B (Grattafiori and others, 2024). This layer—31 of 80—was chosen via cross-validation using an early probe design (see Section˜A.1). Additional experiments on smaller models, such as Llama-3.1-8B and Gemma-3-27B, can be found in Section˜B.7. As training probes is resource-efficient and labels of reasonable quality can realistically be generated using GPT-4o or similar models, we also study how much training on deployment samples (from the development splits, in our case) can help with better capturing the monitored concept.
@@ 5 Related Work
Linear probes. Linear probes (Alain and Bengio, 2018) analyze neural network representations and have been used to understand model internals (Belinkov and Glass, 2019; Tenney et al., 2019; Jin et al., 2025). Alternative designs include Linear Artificial Tomography (Zou et al., 2023) and Mass-Means probes (Marks and Tegmark, 2023). Our work demonstrates how probes can act as computationally effective initial filters. Sparse Auto-Encoders (SAEs) (Cunningham et al., 2023; Bricken et al., 2023) offer an alternative approach via unsupervised clustering, but recent work suggests reliability issues for downstream monitoring tasks (Chanin et al., 2024; Thread, 2023; Kantamneni et al., 2025). We therefore focus on activation probes.
@@
Monitoring for safety-relevant properties. Numerous works detect undesirable LLM behavior via finetuning (Inan et al., 2023a; Zeng et al., 2024). Linear probes offer a cheaper alternative for investigating safety-related concepts like truth and honesty (Marks and Tegmark, 2023; Azaria and Mitchell, 2023; Li et al., 2022; Burns et al., 2023; Panickssery et al., 2023), outperforming prompting whether a statement is true on smaller models in-distribution Liu et al. (2023). While some doubt remains about generalization—Levinstein and Herrmann (2024) show probes from Azaria and Mitchell (2023) fail simple generalization tests, and Belinkov (2021) argues for general skepticism—more recent studies demonstrate robust generalization across transforms and distribution shifts (Wagner, 2024; Roger, 2023). Most recently, Goldowsky-Dill et al. (2025) show deception can be reliably detected across diverse datasets. We apply linear probes to the novel setting of high-stakes user interactions, with rigorous benchmarking against LLM monitors.
Several concurrent works explore related monitoring approaches with different technical choices. Tillman and Mossing (2025) use logistic regression and SAE-based probes on GPT-4o, finding that some probes outperform prompting. This contrasts with our finding that Llama-3.3-70B probes require deployment samples to exceed prompting baselines. We attribute this to: (1) our synthetic-to-real-world OOD evaluation versus their real-data training with smaller distributional shifts, and (2) our extensively optimized prompting baseline (Section˜A.2). Cunningham et al. (2025) highlight the cost-effectiveness of reusing representations for monitoring using either linear probes or partially fine-tuned models, and also explore such monitors as an initial stage in a hierarchical setup. While our work focuses on high-stakes scenarios, theirs centers on jailbreak and harmfulness detection, with many findings aligning with and supporting our own. Oldfield et al. (2025) introduce Truncated Polynomial Classifiers, which augment linear probes with higher-order feature interactions to allow a small amount of dynamic compute scaling within a single activation monitor model.
Decision-making in High-Stakes situations. Human responses under high-stakes conditions have been extensively studied in psychology. Kunreuther et al. (2002) shows how high-stakes scenarios lead to suboptimal decision making, extended by Sosnowski and Brosnan (2023) to nonhuman species. In AI settings, Ren et al. (2025) demonstrates that pressured prompts incentivize LLMs toward dishonesty, while Motwani et al. (2024) shows high-stakes contexts can induce unintended collusion and illicit behaviors like insider trading.